OpenStudy (anonymous):

Medal 25% of what is 44.5?

sammixboo (sammixboo):

Let's use the formula again! \(\large \rm \frac{part}{whole}~=~\frac{percent}{100}\) Percent - 25% The whole is 44.5, so plug them in! \(\large \rm \frac{part}{44.5}~=~\frac{25}{100}\) irst multiply 44.5 by 25 What is 44.5 x 25?

OpenStudy (anonymous):

1112.5

sammixboo (sammixboo):

Correct! now we take the 100 and divide it with 1112.5 So what is 1112.5/100?

OpenStudy (anonymous):

11.125

sammixboo (sammixboo):

Correct! That is our part

sammixboo (sammixboo):

So 11.125 is 25% of 44.5

OpenStudy (anonymous):

so the answer is 11.125?

sammixboo (sammixboo):

Yes!

OpenStudy (anonymous):

>25% of what is 44.5? Just reread the question again, I think you are answering the wrong thing here.

TheSmartOne (thesmartone):

It is \[\frac{ 25 }{ 100 } \times x = 44.5\]

TheSmartOne (thesmartone):

@sammixboo You didn't read the question correctly...

TheSmartOne (thesmartone):

So \[x=\frac{ 44.5 \times 100 }{ 25 }\]

OpenStudy (anonymous):

178

sammixboo (sammixboo):

Dang, again! I keep on messing up. Maybe I shouldn't be doing math when I am sick ahaha

sammixboo (sammixboo):

Now I see! Lol

sammixboo (sammixboo):

Thanks for pointing that out

sammixboo (sammixboo):

I am going to stick with the English section

TheSmartOne (thesmartone):

Yes 178 is correct. :)

TheSmartOne (thesmartone):

Well you must not be that bad if you are a mathlete...

sammixboo (sammixboo):

Nah I just been doing lot's of things wrong since I have been sick I mean I was right, but doing the wrong thing ahaha

sammixboo (sammixboo):

I have done this a multiple times since I was sick, but usually I rarely do this

TheSmartOne (thesmartone):

lol.. Take care of yourself..

sammixboo (sammixboo):

Thanks!